# Generalizing Denoising to Non-Equilibrium Structures Improves Equivariant Force Fields

### v2

@@ A.2 SE(3)/E(3)-Equivariant Networks
Equivariant neural networks (Thomas et al., 2018; Kondor et al., 2018; Weiler et al., 2018; Fuchs et al., 2020; Miller et al., 2020; Townshend et al., 2020; Batzner et al., 2022; Jing et al., 2021; Schütt et al., 2021; Satorras et al., 2021; Brandstetter et al., 2022; Thölke & Fabritiis, 2022; Le et al., 2022; Musaelian et al., 2022; Batatia et al., 2022; Liao & Smidt, 2023; Passaro & Zitnick, 2023; Liao et al., 2023) use vector spaces of irreducible representations (irreps) for equivariant irreps features and adopt equivariant operations such as tensor products to achieve equivariance to 3D rotation. Previous works differ in which equivariant operations are used and the combination of those operations. TFN (Thomas et al., 2018) and NequIP (Batzner et al., 2022) use tensor products for equivariant graph convolution with linear messages, with the latter utilizing extra gate activation (Weiler et al., 2018). SEGNN (Brandstetter et al., 2022) applies gate activation to messages passing for non-linear messages (Gilmer et al., 2017; Sanchez-Gonzalez et al., 2020). SE(3)-Transformer (Fuchs et al., 2020) adopts equivariant dot product attention with linear messages. Equiformer (Liao & Smidt, 2023) improves upon previous models by combining MLP attention and non-linear messages and additionally introducing equivariant layer normalization and regularizations like dropout (Srivastava et al., 2014) and stochastic depth (Huang et al., 2016). However, these networks rely on compute-intensive $SO(3)$ convolutions built from tensor products, and therefore they can only use small values for maximum degrees $L_{max}$ of irreps features. eSCN (Passaro & Zitnick, 2023) significantly reduces the complexity of $SO(3)$ convolutions by first rotating irreps features based on relative positions and then applying $SO(2)$ linear layers, enabling higher values of $L_{max}$. EquiformerV2 (Liao et al., 2023) adopts eSCN convolutions and proposes an improved version of Equiformer to better leverage the power of higher $L_{max}$, achieving the current state-of-the-art results on OC20 (Chanussot* et al., 2021) and OC22 (Tran* et al., 2022) datasets.
We refer readers to the works (Liao & Smidt, 2023; Liao et al., 2023) for detailed background on equivariant networks.
@@ B.1 Training Details
We add an additional block of equivariant graph attention to EquiformerV2 for noise prediction. We mainly follow the hyper-parameters of training EquiformerV2 without DeNS on OC20 S2EF-2M and S2EF-All+MD datasets. For training EquiformerV2 on OC20 S2EF-All+MD dataset, we increase the number of epochs from $1$ to $2$ for better performance. This results in higher training time than other methods. However, we note that we already demonstrate training with DeNS can achieve better results given the same amount of training time in Table 1(a). Table 7 summarizes the hyper-parameters of training EquiformerV2 with DeNS for the ablation studies on OC20 S2EF-2M dataset in Section 4.1.1 and for the main results on OC20 S2EF-All+MD dataset in Section 4.1.2. Please refer to the work of EquiformerV2 (Liao et al., 2023) for details of the architecture. For training eSCN with DeNS, we use the same DeNS-related hyper-parameters as those for training EquiformerV2 for $20$ epochs and the same module as force prediction to predict noise.
@@ C.1 Training Details
Different from OC20, all the atoms in a structure in OC22 are free, and we apply DeNS to all the free atoms. We add an additional block of equivariant graph attention to EquiformerV2 for noise prediction. We follow the same hyper-parameters not relevant to DeNS, and Table 8 summarizes the hyper-parameters for the results on OC22 in Table 3. We use $32$ V100 GPUs (32GB) for training. The training time is $5082$ GPU-hours, and the number of parameters $127$M.
@@ D.2 Training Details
We use the official implementation of Equiformer (Liao & Smidt, 2023) for experiments on the MD17 dataset and follow most of the original hyper-parameters for training with DeNS. For training DeNS, we use an additional block of equivariant graph attention for noise prediction, which slightly increases training time and the number of parameters. The hyper-parameters introduced by training DeNS and the values of energy coefficient $\lambda_{E}$ and force coefficient $\lambda_{F}$ on different molecules can be found in Table 9. Empirically, we find that linearly decaying DeNS coefficient $\lambda_{\text{DeNS}}$ to 00 thoughout the training can result in better performance. For the Equiformer variant without attention and layer normalization, we find that using normal distributions to initialize weights can result in training divergence and therefore we use uniform distributions. For some molecules, we find training Equiformer variant without attention and layer normalization with DeNS is unstable and therefore reduce the learning rate to $3\times 10^{-4}$.
@@
Appendix E Pseudocode for Training with DeNS
@@
Appendix F Visualization of Corrupted Structures

### v3

@@ A.2 SE(3)/E(3)-Equivariant Networks
We first discuss the concept of equivariance and how equivariant networks achieve equivariance and then compare previous works below. We note that most of the content is adapted from Equiformer (Liao & Smidt, 2023) and EquiformerV2 (Liao et al., 2023). We refer readers to the works (Liao & Smidt, 2023; Liao et al., 2023) for more detailed background on equivariant networks and the work (Duval et al., 2024) for a broader review on geometric GNNs for modeling 3D atomistic systems.
3D atomistic systems are often described in 3D coordinate systems. We have the freedom to choose arbitrary 3D coordinate systems since we can change between different coordinates via the symmetries of 3D space. The relevant 3D symmetris are rotation, translation and inversion. The Euclidean group $E(3)$ consists of 3D rotation, translation and inversion while the special Euclidean group $SE(3)$ is comprised of 3D rotation and translation. The laws of physics remain the same regardless of the coordinate we use, and thus the properties of 3D atomistic systems are equivariant to 3D symmetries. For instance, when a 3D atomistic system is rotated, quantities like energy will remain identical while others like forces will rotate accordingly. Mathematically, a function $f$ mapping between vector spaces $X$ and $Y$ is equivariant to a group of transformations $G$ if for any input $x\in X$, output $y\in Y$ and group element $g\in G$, we have $f(D_{X}(g)x)=D_{Y}(g)f(x)=D_{Y}(g)y$, where $D_{X}(g)$ and $D_{Y}(g)$ are transformation matrices or group representations parametrized by $g$ in $X$ and $Y$. Additionally, $f$ is invariant when $D_{Y}(g)$ is an identity matrix for any $g\in G$.
Incorporating equivariance into neural networks as inductive biases can improve data efficiency and generalizability. Equivariant neural networks (Thomas et al., 2018; Kondor et al., 2018; Weiler et al., 2018; Fuchs et al., 2020; Miller et al., 2020; Townshend et al., 2020; Batzner et al., 2022; Jing et al., 2021; Schütt et al., 2021; Satorras et al., 2021; Brandstetter et al., 2022; Thölke & Fabritiis, 2022; Le et al., 2022; Musaelian et al., 2022; Batatia et al., 2022; Liao & Smidt, 2023; Passaro & Zitnick, 2023; Liao et al., 2023) achieve equivariance to 3D rotation and optionally inversion by using vector spaces of irreducible representations (irreps) as equivariant features. The vector spaces of irreps are $(2L+1)$-dimensional, with degree $L$ being a non-negative integer. $L$ can be viewed as the angular frequency of vectors and determines how fast vectors change with respect to a rotation of the coordinate system. Vectors of degree $L$ are referred to as type-$L$ vectors. They are transformed with Wigner-D matrices $D^{(L)}$ when rotating the coordinate system, and $D^{(L)}$ of different $L$ acts on independent vector spaces. Euclidean vectors $\vec{r}$ in $\mathbb{R}^{3}$ such as relative positions and forces can be projected into type-$L$ vectors by using spherical harmonics $Y^{(L)}(\frac{\vec{r}}{||\vec{r}||})$. We concatenate multiple type-$L$ vectors to build an equivariant feature $f$. Given the maximum degree $L_{max}$ of equivariant features, $f$ has $C_{L}$ type-$L$ vectors, where $0\leq L\leq L_{max}$ and $C_{L}$ is the number of channels for type-$L$ vectors. Both $L_{max}$ and $C_{L}$ are architectural hyper-parameters of equivariant networks. Equivariant operations are applied to equivariant features to preserve equivariance, and two examples are tensor products and $SO(3)$ linear operations. Tensor products are the fundamental operation in equivariant networks for interacting vectors of different $L$ and are used to build convolutions and attention. On the other hand, $SO(3)$ linear operations apply separate linear operations to each group of type-$L$ vectors. Relevant to the proposed method, we encode forces into equivariant features by first projecting forces to type-$L$ vectors with spherical harmonics and then expanding the number of channels to $C_{L}$ with an $SO(3)$ linear operation.
@@
Previous works on equivariant networks mainly differ in which equivariant operations are used and the combination of those operations. TFN (Thomas et al., 2018) and NequIP (Batzner et al., 2022) use tensor products for equivariant graph convolution with linear messages, with the latter utilizing extra gate activation (Weiler et al., 2018). SEGNN (Brandstetter et al., 2022) applies gate activation to messages passing for non-linear messages (Gilmer et al., 2017; Sanchez-Gonzalez et al., 2020). SE(3)-Transformer (Fuchs et al., 2020) adopts equivariant dot product attention with linear messages. Equiformer (Liao & Smidt, 2023) improves upon previous models by combining MLP attention and non-linear messages and additionally introducing equivariant layer normalization and regularizations like dropout (Srivastava et al., 2014) and stochastic depth (Huang et al., 2016). However, these networks rely on compute-intensive $SO(3)$ convolutions built from tensor products, and therefore they can only use small values for maximum degrees $L_{max}$ of irreps features. eSCN (Passaro & Zitnick, 2023) significantly reduces the complexity of $SO(3)$ convolutions by first rotating irreps features based on relative positions and then applying $SO(2)$ linear layers, enabling higher values of $L_{max}$. EquiformerV2 (Liao et al., 2023) adopts eSCN convolutions and proposes an improved version of Equiformer to better leverage the power of higher $L_{max}$, achieving the current state-of-the-art results on OC20 (Chanussot* et al., 2021) and OC22 (Tran* et al., 2022) datasets.
A.3 Equiformer Series
Equiformer (Liao & Smidt, 2023) is an $SE(3)$/$E(3)$-equivariant graph neural network that combines the inductive biases of 3D-related equivariance with the strength of Transformers (Vaswani et al., 2017). Starting from Transformers, Equiformer introduces three architectural modifications. First, Equiformer adopts equivariant features built from vector spaces of irreps as internal representations to incorporate equivariance. Second, equivariant operations are applied to the equivariant features. These operations include tensor products and the equivariant counterparts of the original operations in Transformers. The latter part consists of equivariant linear operations (i.e., $SO(3)$ linear operations), equivariant layer normalization (Ba et al., 2016) and gate activation (Weiler et al., 2018). Third, Equiformer proposes to apply non-linear functions to both attention weights and message passing, which improves the expressivity of attention in standard Transformers.
Although Equiformer demonstrates that Transformers generalize well to 3D atomistic systems, it is limited to small values of maximum degree $L_{max}$ because of the compute-intensive tensor product operations. Higher degrees can better capture angular resolutions and directional information, and therefore lower $L_{max}$ can limit the expressivity of Equiformer. To address this limitation, EquiformerV2 (Liao et al., 2023) adopts eSCN (Passaro & Zitnick, 2023) convolutions, which significantly reduce the computational complexity of tensor products, to incorporate higher-degree equivariant representations and proposes three architectural improvements to better leverage the power of higher degrees. First, attention re-normalization is proposed to introduce one additional layer normalization to the non-linear functions of attention weights. This helps stabilize attention and improves empirical performance. Second, separable $S^{2}$ activation based on $S^{2}$ activation (Cohen et al., 2018) is proposed to better mix the information of all degrees and stabilize training. Third, separable layer normalization is proposed to replace the original equivariant layer normalization in order to preserve the relative importance of different degrees.
@@ B.1 Training Details
For force prediction, we adopt direct methods following previous works for a fair comparison. We add an additional block of equivariant graph attention to EquiformerV2 for noise prediction. We mainly follow the hyper-parameters of training EquiformerV2 without DeNS on OC20 S2EF-2M and S2EF-All+MD datasets. For training EquiformerV2 on OC20 S2EF-All+MD dataset, we increase the number of epochs from $1$ to $2$ for better performance. This results in higher training time than other methods. However, we note that we already demonstrate training with DeNS can achieve better results given the same amount of training time in Table 1(a). Table 7 summarizes the hyper-parameters of training EquiformerV2 with DeNS for the ablation studies on OC20 S2EF-2M dataset in Section 4.1.1 and for the main results on OC20 S2EF-All+MD dataset in Section 4.1.2. Please refer to the work of EquiformerV2 (Liao et al., 2023) for details of the architecture. For training eSCN with DeNS, we use the same DeNS-related hyper-parameters as those for training EquiformerV2 for $20$ epochs and the same module as force prediction to predict noise.
@@ C.1 Training Details
Different from OC20, all the atoms in a structure in OC22 are free, and we apply DeNS to all the free atoms. We add an additional block of equivariant graph attention to EquiformerV2 for noise prediction. We follow the same practice as on OC20 and use direct methods for force prediction. We follow the same hyper-parameters not relevant to DeNS, and Table 8 summarizes the hyper-parameters for the results on OC22 in Table 3. We use $32$ V100 GPUs (32GB) for training. The training time is $5082$ GPU-hours, and the number of parameters is $127$M.
@@ D.2 Training Details
We use the official implementation of Equiformer (Liao & Smidt, 2023) for experiments on the MD17 dataset and follow most of the original hyper-parameters for training with DeNS. Gradient methods are used to predict forces. For training DeNS, we use an additional block of equivariant graph attention for noise prediction, which slightly increases training time and the number of parameters. The hyper-parameters introduced by training DeNS and the values of energy coefficient $\lambda_{E}$ and force coefficient $\lambda_{F}$ on different molecules can be found in Table 9. Empirically, we find that linearly decaying DeNS coefficient $\lambda_{\text{DeNS}}$ to 00 thoughout the training can result in better performance. For the Equiformer variant without attention and layer normalization, we find that using normal distributions to initialize weights can result in training divergence and therefore we use uniform distributions. For some molecules, we find training Equiformer variant without attention and layer normalization with DeNS is unstable and therefore reduce the learning rate to $3\times 10^{-4}$.
@@
D.3 Additional Simulation-Based Results
Following the work (Fu et al., 2023), we run simulations on the four molecules (i.e., Aspirin, Ethanol, Naphthalene and Salicylic Acid) and compare the two simulation-based metrics, which are stability and distribution of interatomic distances $h(r)$. We use the previously trained Equiformer for the simulations. We note that Equiformer models are trained on $950$ examples for each molecule instead of $9,500$ as in Fu et al. (2023). The results are summarized in Table 10. Training with DeNS helps energy and forces MAE as well as stability. The results of $h(r)$ are similar. For Naphthalene, training with DeNS improves stability from $133.8/300$ to $157.2/300$. Both models become unstable when running simulations of Naphthalene, and we surmise that is because we only use $950$ examples for training instead of $9,500$. For others, the performance gain in simulation-based metrics is not significant since the original Equiformer trained on $950$ examples already achieves similar results to other top-performing models trained on $9,500$ examples, suggesting that the potential room for improvement would be quite limited.
Aspirin Ethanol Naphthalene Salicylic acid
Model $\downarrow$ $\downarrow$ $\uparrow$ $h(r)\downarrow$ $\downarrow$ $\downarrow$ $\uparrow$ $h(r)\downarrow$ $\downarrow$ $\downarrow$ $\uparrow$ $h(r)\downarrow$ $\downarrow$ $\downarrow$ $\uparrow$ $h(r)\downarrow$
Equiformer ($L_{max}=2$) 5.3 7.2 300 0.02 2.2 3.1 289.9 0.09 3.7 2.1 133.8 0.12 4.5 4.1 300 0.03
Equiformer ($L_{max}=2$) + DeNS 5.1 5.7 300 0.02 2.2 2.6 300 0.09 3.7 1.7 157.2 0.12 4.4 3.7 300 0.03
Table 10: Simulation-based results on MD17 dataset. We report simulation-based metrics, stability and distribution of interatomic distances $h(r)$. Models are the same as in Table 4. Energy and force are in units of meV and meV/Å and are evaluated on the testing set.
Appendix E Pseudocode for Force Encoding
We provide the pseudocode for force encoding in Algorithm 1. Here the original node embedding $x_{i}$ contains only the atom embedding $x_{i,z}$. Note that the type-$L$ vectors in $x_{i,z}$ are all zeros for $L>0$ since $x_{i,z}$ is obtained by applying an $SO(3)$ linear layer to type-00 vectors. We directly add the force embedding $x_{i,\mathbf{f}}$ to the original node embedding to encode forces.
Algorithm 1 Force Encoding
1:$x_{i,z}\leftarrow\text{SO3\_Linear}\left(\text{one\_hot}(z_{i})\right)$triangleright Extend Equation 5 to all degrees $L$ and apply to the one-hot encoding of atomic numbers $z_{i}$ for each atom
2:$x_{i}\leftarrow x_{i,z}$
3:$x_{i,\mathbf{f}}\leftarrow\text{SO3\_Linear}\left(||\mathbf{f}_{i}||\cdot Y\left(\frac{\mathbf{f}_{i}}{||\mathbf{f}_{i}||}\right)\right)$
4:$x_{i}\leftarrow x_{i}+x_{i,\mathbf{f}}$
Appendix F Pseudocode for Training with DeNS
@@
Appendix G Visualization of Corrupted Structures